# Mixing Natural and Synthetic Images for Robust Self-Supervised Representations

### v1

@@ Abstract
This paper introduces DiffMix, a new self-supervised learning (SSL) pre-training framework that combines real and synthetic images. Unlike traditional SSL methods that predominantly use real images, DiffMix uses a variant of Stable Diffusion to replace an augmented instance of a real image, facilitating the learning of cross real-synthetic image representations. The key insight is that while SSL methods trained solely on synthetic images underperform compared to those trained on real images, a blended training approach using both real and synthetic images leads to more robust and adaptable representations. Experiments demonstrate that DiffMix enhances the SSL methods SimCLR, BarlowTwins, and DINO, across various robustness datasets and domain transfer tasks. DiffMix boosts SimCLR’s accuracy on ImageNet-1K by 4.56%. These results challenge the notion that high-quality real images are crucial for SSL pre-training by showing that lower quality synthetic images can also produce strong representations. DiffMix also reduces the need for image augmentations in SSL, offering new optimization strategies.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Comparison of our proposed DiffMix pipeline against existing Joint embedding/Distillation-based SSL methods that use a Siamese setting. DiffMix modifies a branch of augmentations to replace real image $x^{\prime}_{1}$ with a synthetically generated image $\tilde{x}_{1}$
@@
There has been a growing interest in using synthetic images for supervision (Sariyildiz et al., 2022; He et al., 2022b, a; Tian et al., 2023). These methods involve creating artificial images using generative models like GANs (Goodfellow et al., 2014) and Diffusion models (Ho et al., 2020), providing a vast and flexible data source. The appeal of synthetic image datasets lies in their ease of generation, wider range of semantic content, and minimal human intervention, addressing key concerns in computer vision like cost efficiency and fairness in data collection and annotation. Although generative models address the issue of data scarcity, exclusive reliance on synthetic images for supervision is not without drawbacks. The main challenge has been the domain gap between synthetic and real-world data. Models trained only on synthetic images often struggle to adapt to real-world settings due to their limited exposure to the variability and complexity of natural images (Tremblay et al., 2018). This issue is particularly pronounced in large-scale image recognition tasks, where models trained on synthetic data typically underperform those trained on real images (Vanherle et al., 2022; Jeelani et al., 2023). In response to this, our research proposes a novel training framework that integrates both real and synthetic images, aiming to harness the strengths of each data source while mitigating their individual weaknesses.
On the issue of annotated data scarcity, self-supervised learning (SSL) has concurrently enjoyed significant advancements in recent years. The capability of unsupervised learning to generate high-quality features using pretext tasks now parallels, and in some cases surpasses, that of supervised learning. Even though SSL saves the annotation costs, the preparation of training data is still a vexing problem. Current research within SSL predominantly concentrates on the development of the pretext tasks, while the characteristics of the data being utilized remain less explored. (Tian et al., 2023) uses text-conditional image generator for the creation of synthetic images, which are subsequently employed in the pre-training of an SSL model. This approach necessitates the use of image captions and a substantial dataset ($\sim 20M$ images) to achieve performance parity with SSL models trained on real data. A training framework that uses synthetic images without any form of labels, prompts or huge amounts of training data, paired with real data has not been explored in the SSL setting before.
We ask the following question: Is it possible to improve SSL representations by combining both real and synthetic images using generative models in a fully self-supervised way? This paper answers this question in the affirmative and provide DiffMix, that explores the potential of combining synthetic images, generated by generative models without any labelled data, with real-world images for SSL training. We illustrate DiffMix in Figure 1. This approach aims to leverage the advantages of both data types: the controlled diversity of synthetic images and the authenticity of real-world data. We find that models pre-trained exclusively on synthetic images underperform compared to those pre-trained with real images across most scenarios. Interestingly, the combined pre-training approach using both real and synthetic images of DiffMix markedly boosts the performance of the same model, indicating superior learned representations. Additionally, DiffMix can work with virtually any model architecture with little modification, indicating that mixed pre-training strategy might be superior to conventional SSL pre-training methods.
We applied the mixing technique of DiffMix to various established SSL methods, including SimCLR(Chen et al., 2020), DINO (Caron et al., 2021), and BarlowTwins (Zbontar et al., 2021), and observed notable improvements in their performance. We complement the results with attributes of the generated synthetic images that we find are pivotal in how effectively the models learn and adapt to new datasets, particularly in aspects of robustness and adaptability. Contrary to the common belief that SSL performs better with high-quality real images, our research shows that lower quality synthetic images may contribute to stronger representations. Interestingly, we also discovered that image augmentations, typically beneficial for SSL, become somewhat unnecessary and even redundant in a mixed real and synthetic image setting. Through an analysis of various synthetic image qualities and their effects on the mixing process, we showcase new insights into optimizing SSL pre-training strategies. Our key contributions are:
  • We introduce DiffMix, a new SSL pre-training framework, which facilitates the learning of cross real-synthetic image representations without any labels.
  • The representations learned through DiffMix are found to be more adaptable than representations learned from real images towards distribution shifts and various domain transfer tasks. We report up to 4.56% improvement on ImageNet-1K using SimCLR.
  • We discover that lower-quality synthetic data leads to better mixing and superior performance. When trained without augmentation, DiffMix even surpasses SimCLR trained with heavy augmentation, suggesting synthetic data may replace augmentations in SSL.
@@ 2 Related Work
While SSL techniques exist in different forms, one of the most successful self-supervised learning paradigms is joint-embedding SSL (Misra & Maaten, 2020; He et al., 2020; Ye et al., 2019; Oord et al., 2018; Chen et al., 2020; Zbontar et al., 2021). The main focus of joint-embedding SSL is instance-based discriminative learning (Dosovitskiy et al., 2014; Assran et al., 2023), where each image is considered to be its own class, and a model is trained by discriminating different views of the same image generated using data augmentation (He et al., 2020; Chen et al., 2020; Zbontar et al., 2021; Caron et al., 2021). One such example is SimCLR (Chen et al., 2020), which uses an InfoNCE-based formulation (Oord et al., 2018) to bring in the representation of different views of the same image closer (‘positive pairs’), and repel representations of views from different images (‘negative pairs’) apart. In most cases, joint-embedding SSL methods work in a Siamese setting(Chen & He, 2021) where two branches have identical architectures and share weights. However, networks under such the Siamese setting are vulnerable to collapsing to trivial representations. BarlowTwins (Zbontar et al., 2021) brings covariance regularization to the contrastive setting as a to enforce a non-collapsing solution. More recently, works such as DINO (Caron et al., 2021) have shown alternative ways to prevent collapse using architectural strategies inspired by knowledge distillation (Hinton et al., 2015). In this work, we provide an improved pre-training mechanism for representation learning in such joint-embedding SSL techniques.
@@
Recent advancements in machine learning have increasingly leveraged synthetic data across a variety of domains (Silver et al., 2017; Mimura et al., 2018; Meng et al., 2022; Dan et al., 2019). This type of data is particularly crucial for tasks that demand extensive labeled datasets, such as human pose estimation (Ma et al., 2022; Guo et al., 2022), semantic segmentation (Chen et al., 2018; Rewatbowornwong et al., 2021), and optical flow estimation (whan Kim et al., 2022; Sun et al., 2021). In the task of image classification, several studies have demonstrated the effectiveness of synthetic data. (He et al., 2022b) illustrates its application in data-scarce settings and transfer learning; (Wang et al., 2023) explores its role in enhancing adversarial training; (Bansal & Grover, 2023; Sariyildiz et al., 2022) evaluates model robustness against natural distribution shifts using synthetic data; and (Azizi et al., 2023) discusses augmentation of images through fine-tuned text-to-image diffusion models. It is important to note that all of these studies focus on supervised learning. Our work, however, is distinct in its concentration on SSL and additionally investigates both robustness and transfer learning. The study most closely aligned with our research is that of (Tian et al., 2023), which utilizes text-to-image diffusion models for generating images, often producing multiple images from a single caption. In contrast, our approach employs image-to-image diffusion models and is designed to generate only one image for each source image. This method reduces the reliance on labeled data for image generation and the cost of image generation.
Generative Models. The landscape of synthetic image generation has seen a significant evolution, with Generative Adversarial Networks (GANs) such as BigGAN (Brock et al., 2018) initially setting a high standard. These models have been pivotal in pushing the boundaries of image realism and quality. Recently, diffusion models have emerged as a promising alternative, demonstrating impressive results in both conditional (Dhariwal & Nichol, 2021; Saharia et al., 2021) and unconditional (Ho et al., 2021) synthetic image generation. Text-to-image diffusion models like DALL-E (Ramesh et al., 2021) and Imagen (Saharia et al., 2022) are notable examples, showcasing the ability to create detailed and contextually accurate images from textual descriptions. Our research takes a unique turn by focusing on the image-to-image diffusion model, specifically a fine-tuned version of Stable Diffusion (Rombach et al., 2021). This model distinguishes itself by utilizing CLIP (Radford et al., 2021) image embeddings instead of text embeddings.
@@ 3 Method
In this work, we primarily focus on Self-Supervised Learning (SSL) techniques, particularly those that consolidate representations from different perspectives or augmentations of the same instance(Alexey et al., 2016; Wu et al., 2018; He et al., 2020; Chen et al., 2020). The main idea behind this technique is, through iterative processes, these representations gradually become less sensitive to the transformations generating these varied views. Consequently, this leads to the learning of image representations that are notably effective for vision tasks such as classification(He et al., 2020; Chen et al., 2020; Zbontar et al., 2021). In this section, we introduce our framework, DiffMix, which uniquely employs both real and synthetically generated data through stable diffusion, and see how it can be incorporated in some of the existing SSL frameworks.
3.1 Description of DiffMix
@@
Consider $x_{1}$ and $x^{\prime}_{1}$, two augmented patches from an image, randomly selected from a dataset. These augmentations can include a variety of changes, such as altering spatial positions within an image, adding varying noise and applying random color adjustments, etc. Existing instance-based discriminative SSL methods primarily rely on real images(He et al., 2020; Chen et al., 2020; Zbontar et al., 2021; Caron et al., 2021). In these methods, the representation derived from the first augmentation, $x_{1}$, of a real image is anticipated to closely align with the representation of the second augmentation, $x_{1}^{\prime}$, of the same image as shown in Figure 1. Our DiffMix framework modifies this approach by incorporating synthetically generated images alongside real ones. The primary objective of DiffMix is to synchronize the representations of real and synthetic images, thus enhancing existing SSL methodologies such as SimCLR, DINO, and BarlowTwins.
@@
The synthetic images employed by DiffMix have a unique characteristic that they share a variation of the same semantic component or object with that of the real image. To achieve this, we employ Image Variation Diffuser [^1] https://huggingface.co/lambdalabs/sd-image-variations-diffusers, a variant of Stable Diffusion (Rombach et al., 2021) tailored to generate diverse images while preserving semantic categories or in simple terms, the image class. In the Stable Diffusion-based generative model, represented as $g^{k}_{SD}(\cdot)$, where $k$ indicates the guidance scale influencing the generative features from the input image, an input $x_{i}\sim D$ yields a synthetic counterpart $\tilde{x_{i}}$, such that $\tilde{x_{i}}=g^{k}_{SD}(x_{i})$ The innovative aspect of DiffMix lies in substituting a portion of the augmentation process, i.e, the second branch of augmentation $x_{i}^{\prime}$ within the SSL framework with these synthetic images $\tilde{x_{i}}$, to learn cross real-synthetic image representations a shown in Figure 1.
@@
Figure 2: Existing SSL methods, including (A) SimCLR, (B) Barlow Twins, and (C) DINO, have been enhanced with our novel DiffMix approach. In both (A) SimCLR and (B) Barlow Twins, we replace a branch representing the positive pair with a synthetic image generated without the label using Stable Diffusion. This modification enables the learning of real-synthetic view prediction. (C) DINO utilizes a distillation framework with two global views for the teacher and a mix of two global and eight local views for the student. Our adaptation integrates an equal blend of global and local synthetic and real images facilitating learning correspondences between global-to-local on top of real-to-synthetic image views.
@@ 3.2 Mixing in joint-embedding SSL
SimCLR + DiffMix:
In SimCLR’s contrastive learning framework (Chen et al., 2020), ’positive‘ and ’negative‘ pairs of images are identified, with the goal of either converging or diverging their representations. In SimCLR, two augmented views are generated for each image in a mini-batch, resulting in $2N$ images for a mini-batch size of $N$. Each view is paired with its corresponding alternate view as a ’positive‘ pair, while the remaining $2(N-1)$ images are treated as ’negative’ pairs. We now propose to incorporate mixing into SimCLR. We first define a new set $\{x_{k},\tilde{x}_{k}\}$ for $k\in[1,2,...,N]$, where $\tilde{x}_{k}$ denotes the synthetically generated counterpart of $x_{k}$, thus establishing pairs like $x_{1}$ and $\tilde{x}_{1}$ as positive examples. The contrastive prediction task of the modified, which we term as SimCLR+DiffMix now involves identifying $x_{1}$ and $\tilde{x}_{1}$ in $\{x_{k},\tilde{x}_{k}\}\,\forall\,k\in[1,N]$. The modified loss function for the mixed version of SimCLR, denoted as $\mathcal{L}_{MixSR}$, for a positive pair of examples ($x_{i}$, $\tilde{x_{i}}$), is defined as:
@@
While the original SimCLR focuses on aligning representations of variously distorted versions of the same real image, our modified approach aims to align representations of a real image with its synthetic counterpart generated by Stable Diffusion as shown in Figure 2 (A). Concurrently, we strive to differentiate these pairs from other real and synthetic image pairs.
Barlow Twins + DiffMix:
The Barlow Twins framework (Zbontar et al., 2021), while maintaining a Siamese network structure similar to SimCLR (Chen et al., 2020), adopts a distinct approach to representation alignment. The difference lies in the Barlow Twins’ objective function, which assesses the cross-correlation matrix between the embeddings from two identical networks. These networks process distorted versions of a batch of samples, with the aim of aligning this matrix closely with the identity matrix. This alignment ensures that the embeddings of distorted versions of a sample are similar, while simultaneously reducing redundancy among the components of these embeddings.

### v2

# MixDiff: Mixing Natural and Synthetic Images for Robust Self-Supervised Representations

@@ Abstract
This paper introduces MixDiff, a new self-supervised learning (SSL) pre-training framework that combines real and synthetic images. Unlike traditional SSL methods that predominantly use real images, MixDiff uses a variant of Stable Diffusion to replace an augmented instance of a real image, facilitating the learning of cross real-synthetic image representations. Our key insight is that while models trained solely on synthetic images underperform, combining real and synthetic data leads to more robust and adaptable representations. Experiments show MixDiff enhances SimCLR, BarlowTwins, and DINO across various robustness datasets and domain transfer tasks, boosting SimCLR’s ImageNet-1K accuracy by 4.56%. Our framework also demonstrates comparable performance without needing any augmentations, a surprising finding in SSL where augmentations are typically crucial. Furthermore, MixDiff achieves similar results to SimCLR while requiring less real data, highlighting its efficiency in representation learning [^1] We have made the source code and generated data available to the public at: https://github.com/cryptonymous9/mixing-ssl.
@@ 1 Introduction
Figure 1: Comparison of SimCLR performance on real, synthetic (Syn), and mixed real and synthetic images (MixDiff). The radar charts show normalized accuracy across 8 transfer learning datasets (left) and ImageNet-1K plus 6 distribution shift datasets (right), with values from 0.5 to 1.1. MixDiff enhances in-distribution and robustness performance and generalizes better. More details in Sec. 4.
Self-supervised learning (SSL) has enjoyed significant advancements in recent years [26]. The capability of joint-embedding SSL architectures to generate high-quality features using pretext tasks now parallels, and in some cases surpasses, that of supervised learning. Joint-embedding SSL methods can include distillation [76, 12, 25] or contrastive strategies [38, 13, 29, 81], where multiple network branches aim to learn representations by maximizing agreement between differently augmented views of the same data example in the embedding space. Even though such recent advancements in SSL save annotation costs, preparing training data is still challenging. Current research within SSL predominantly concentrates on the development of the pretext tasks, while the characteristics of the data being utilized for learning remain less explored.
@@
This focus on pretext tasks, however, has led researchers to consider alternative data sources to further improve model performance. In particular, there has been a growing interest in using synthetic images for self-supervision [74, 62, 70]. The appeal of synthetic image datasets lies in their ease of generation, a wider range of semantic content, and minimal human intervention, addressing key concerns in computer vision like cost efficiency and fairness in data collection and annotation. Although generative models address the issue of data scarcity, exclusive reliance on synthetic images for supervision is not without drawbacks. The main challenge has been the domain gap between synthetic and real-world data. Models trained only on synthetic images often struggle to adapt to real-world settings due to their limited exposure to the variability and complexity of natural images [66]. This issue is particularly pronounced in large-scale image recognition tasks, where models trained on synthetic data typically underperform those trained on real images [67, 36]. In response to this, our research proposes a novel training framework that integrates both real and synthetic images, aiming to harness the strengths of each data source while mitigating their individual weaknesses.
In this paper, we introduce MixDiff to explore the potential of combining synthetic images generated by generative models without any labeled data with real-world images for SSL training. MixDiff is a simple framework that replaces an augmented instance of a real image in an existing joint-embedding SSL pipeline with a synthetic image from a generative model. The simplicity of the framework allows it to incorporate it in existing SSL methods like SimCLR [13], BarlowTwins [81] and DINO [12] as shown in Figure 2.
MixDiff works on the idea that synthetically generated images generate harder positive pairs that makes the overall training objective less trivial [55]. We find that SSL models pre-trained exclusively on synthetic images underperform compared to those pre-trained with real images across most scenarios. Interestingly, our proposed MixDiff, which uses both real and synthetic images, improves model performance in SSL not only on in-distribution datasets but also on various out-of-distribution tasks as we show in Figure 1, suggesting enhanced representation learning. Specifically, we observe an average increase in top-1 accuracy of about 26.92% across six distributional datasets and a 7.36% improvement in transfer learning across eight datasets. This observation suggests that while synthetic images alone may be insufficient for optimal pre-training, MixDiff capitalizes on the synergistic approach of leveraging the strengths of both image types to learn more robust SSL representations.
Building on these results, we investigated how synthetic image quality affects model performance in SSL. Prior research has shown a strong correlation between the quality of diffusion-generated synthetic images and model performance. Our study reveals that MixDiff is less sensitive to variations in synthetic image quality, potentially reducing the need for precise quality optimization. We also find that integrating synthetic data in models like DINO may de-emphasize background features, suggesting enhanced scene layout understanding beneficial for image segmentation. Despite the computational cost of generating synthetic images, MixDiff requires fewer real images to match SimCLR performance, indicating more efficient SSL pre-training. The reduced reliance on large real datasets and high-quality synthetic images, coupled with robustness to distributional shifts and transfer learning, positions MixDiff as a promising approach for enhancing SSL pre-training.
@@ 2 Related Work
While SSL techniques exist in different forms, one of the most successful self-supervised learning paradigms is joint-embedding SSL [46, 29, 78, 48, 13, 81]. The main focus of joint-embedding SSL is instance-based discriminative learning [22, 3], where each image is considered to be its own class, and a model is trained by discriminating different views of the same image generated using data augmentation [29, 13, 81, 12]. One such example is SimCLR [13], which uses an InfoNCE-based formulation [48] to bring in the representation of different views of the same image closer (positive pairs) and repel representations of views from different images (negative pairs) apart. In most cases, joint-embedding SSL methods work in a Siamese setting[14] where two branches have identical architectures and share weights. However, networks such as the Siamese setting are vulnerable to collapsing to trivial representations. BarlowTwins [81] brings covariance regularization to the contrastive setting to enforce a non-collapsing solution. More recently, works such as DINO [12] have shown alternative ways to prevent collapse using architectural strategies inspired by knowledge distillation [34] and addressing catastrophic forgetting [6]. We provide an improved pre-training mechanism for representation learning in such joint-embedding SSL techniques.
@@
Recent advancements in machine learning have increasingly leveraged synthetic data across a variety of domains [60, 45, 44, 28]. This type of data is particularly crucial for tasks that demand extensive labeled datasets, such as human pose estimation [42, 27], semantic segmentation [15, 54], optical flow estimation [73, 61], and language models [1, 24, 68]. In the task of image classification, several studies have demonstrated the effectiveness of synthetic data [72]. [30, 65] illustrates its application in data-scarce settings and transfer learning; [71] explores its role in enhancing adversarial training; [23] diversifies images; [7, 59] evaluates model robustness against natural distribution shifts using synthetic data; and [65, 4] discusses augmentation of images through fine-tuned diffusion models. It is important to note that all of these studies focus on supervised learning. Our work, however, is distinct in its concentration on SSL. Recently, there has been significant interest in leveraging synthetic data for SSL [74, 62, 70]. [62] employ text-to-image diffusion models to generate multiple images from a single caption, while our approach uses image-to-image diffusion models to produce one image per source, reducing the dependency on labeled data and associated costs. [74] introduces a data generation framework to enhance contrastive learning, with the generator trained with the SSL model. [70] train diffusion models and mix real and synthetic data to boost contrastive learning through data augmentation and inflation. Unlike these methods, our approach does not require training a new generative model. Instead, we utilize off-the-shelf variants of Stable Diffusion[56, 77] to improve the quality of representations in existing SSL models.
Generative Models.
The landscape of synthetic image generation has seen a significant evolution, with Generative Adversarial Networks (GANs) such as BigGAN [11] initially setting a high standard. These models have been pivotal in pushing the boundaries of image realism and quality. Recently, diffusion models have emerged as a promising alternative, demonstrating impressive results in both conditional [19, 58] and unconditional [35] synthetic image generation. Text-to-image diffusion models like DALL-E [52] and Imagen [57] are notable examples, showcasing the ability to create detailed and contextually accurate images from textual descriptions. Our research takes a unique turn by focusing on the image-to-image diffusion model, specifically a fine-tuned version of Stable Diffusion [56]. This model distinguishes itself by utilizing CLIP [51] image embeddings instead of text embeddings.
@@ 3 Method
In this work, we focus on Self-Supervised Learning (SSL) techniques, particularly those that consolidate representations from different perspectives or augmentations of the same instance [2, 75, 29, 13]. The main idea behind this technique is, through iterative processes, these representations gradually become less sensitive to the transformations generating these varied views. Consequently, this leads to the learning of image representations that are notably effective for vision tasks such as classification [29, 13, 81]. In this section, we introduce our framework, MixDiff, which uniquely employs both real and synthetically generated data through stable diffusion, and see how it can be incorporated in some of the existing SSL frameworks.
3.1 Description of MixDiff
@@
Consider $x_{1}$ and $x^{\prime}_{1}$, two augmented patches from an image, randomly selected from a dataset. These augmentations can include a variety of changes, such as altering spatial positions within an image, adding varying noise and applying random color adjustments, etc. Existing instance-based discriminative SSL methods primarily rely on real images [29, 13, 81, 12]. In these methods, the representation derived from the first augmentation, $x_{1}$, of a real image is anticipated to closely align with the representation of the second augmentation, $x_{1}^{\prime}$, of the same image as shown in Figure 1. Our MixDiff framework modifies this approach by incorporating synthetically generated images alongside real ones. The primary objective of MixDiff is to synchronize the representations of real and synthetic images, thus enhancing existing SSL methodologies such as SimCLR, DINO, and BarlowTwins.
@@
The synthetic images employed by MixDiff have a unique characteristic that they share a variation of the same semantic component or object with that of the real image. To achieve this, we employ Image Variation Diffuser  [50] [^2] https://huggingface.co/lambdalabs/sd-image-variations-diffusers, a variant of Stable Diffusion (SD) [56] tailored to generate diverse images while preserving semantic categories or in simple terms, the image class. In the SD-based generative model, represented as $g^{k}_{SD}(\cdot)$, where $k$ indicates the guidance scale influencing the generative features from the input image, an input $x_{i}\sim D$ yields a synthetic counterpart $\tilde{x_{i}}$, such that $\tilde{x_{i}}=g^{k}_{SD}(x_{i})$ The innovative aspect of MixDiff lies in substituting a portion of the augmentation process, i.e, the second branch of augmentation $x_{i}^{\prime}$ within the SSL framework with these synthetic images $\tilde{x_{i}}$, to learn cross real-synthetic image representations a shown in Figure 1.
@@
Figure 2: Existing SSL methods, including (A) SimCLR, (B) Barlow Twins, and (C) DINO, have been enhanced with our novel MixDiff approach. In both (A) SimCLR and (B) Barlow Twins, we replace a branch representing the positive pair with a synthetic image generated without the label using Stable Diffusion. This modification enables the learning of real-synthetic view prediction. (C) DINO utilizes a distillation framework with two global views for the teacher and a mix of two global and eight local views for the student. Our adaptation integrates a blend of global and local synthetic and real images facilitating learning correspondences between global-to-local on top of real-to-synthetic image views.
@@ 3.2 Mixing in joint-embedding SSL
SimCLR + MixDiff:
In SimCLR’s contrastive setup [13], ’positive‘ and ’negative‘ pairs of images are identified, with the goal of either converging or diverging their representations. In SimCLR, two augmented views are generated for each image in a mini-batch, resulting in $2N$ images for a mini-batch size of $N$. Each view is paired with its corresponding alternate view as a ’positive‘ pair, while the remaining $2(N-1)$ images are treated as ’negative’ pairs. We now propose to incorporate mixing into SimCLR. We first define a new set $\{x_{k},\tilde{x}_{k}\}$ for $k\in[1,2,...,N]$, where $\tilde{x}_{k}$ denotes the synthetically generated counterpart of $x_{k}$, thus establishing pairs like $x_{1}$ and $\tilde{x}_{1}$ as positive examples. The contrastive prediction task of the modified, which we term as SimCLR+MixDiff now involves identifying $x_{1}$ and $\tilde{x}_{1}$ in $\{x_{k},\tilde{x}_{k}\}\,\forall\,k\in[1,N]$. The modified loss function for the mixed version of SimCLR, denoted as $\mathcal{L}_{MixSR}$, for a positive pair of examples ($x_{i}$, $\tilde{x_{i}}$), is defined as:
@@
Barlow Twins + MixDiff:
The Barlow Twins framework [81], while maintaining a Siamese network structure similar to SimCLR [13], adopts a distinct approach to representation alignment. The difference lies in the Barlow Twins’ objective function, which assesses the cross-correlation matrix between the embeddings from two identical networks. These networks process distorted versions of a batch of samples, with the aim of aligning this matrix closely with the identity matrix. This alignment ensures that the embeddings of distorted versions of a sample are similar, while simultaneously reducing redundancy among the components of these embeddings.